OpenStudy (anonymous):

\(\sqrt{125b}\) Simplify

OpenStudy (anonymous1):

Factor 125: \[125=5^3=5^2*5\] \[\sqrt{125b}=\sqrt{5^2*5b}=\sqrt{5^2}*\sqrt{5b} \\=5\sqrt{5b}\]